### v2

# Intent-Based Network for RAN Management with Large Language Models

Abstract
Advanced intelligent automation becomes an important feature to deal with the increased complexity in managing wireless networks. This paper proposes a novel intent-based automation for Radio Access Network (RAN) management, leveraging Large Language Models (LLMs). The proposed method enhances intent translation by autonomously interpreting high-level objectives, reasoning over complex network states, and generating precise RAN configurations through an LLM-integrated agentic architecture. We propose a structured prompt engineering technique and demonstrate that the network can automatically improve its energy efficiency by dynamically optimizing critical RAN parameters through a closed-loop mechanism. It showcases the potential for robust resource management in RAN by adapting strategies based on real-time feedback via LLM-orchestrated agentic systems.
Index Terms:
intent-based network, agentic AI, LLM, Wireless Network
I Introduction
The rapid increase in diverse service types and dynamic resource demands in RAN introduces unprecedented operational complexity. Even though Open RAN enables the integration of components from multiple vendors, it can also increase configuration complexity and the risks of misconfiguration. This configuration heavily relies on low-level, manual procedures, which makes agility and scalability difficult to achieve [1].
Intent-Based Network (IBN) has emerged to manage networks by simplifying detailed technical configurations with minimal external intervention [2]. An intent serves as a high-level, human-expressible declaration of network goals, rather than specified methods for achievement. It acts as a structured set of expectations, encompassing requirements, goals, conditions, and constraints [3]. An IBN then implements and manages these intents [4], fundamentally abstracting network complexity and allowing users and customers to request services without requiring detailed knowledge of their underlying provision. It is crucial to note that "intent" is a fundamental term with various interpretations across key Internet and Telco bodies. The Internet Research Task Force (IRTF) in RFC 9315 [4] defines intent as a high-level, declarative specification of network goals and desired outcomes, emphasizing it should be network-wide, outcome-driven, and vendor-agnostic. Similarly, 3GPP, in its standardization for 5G networks, defines intent as a set of expectations including requirements, goals, conditions, and constraints, focusing on "what" needs to be achieved [3]. They further categorize intents based on roles such as Intent-CSC, Intent-CSP, and Intent-NOP. TM Forum’s IG1230 document [5] concurs, defining Intent as "the formal specification of the expectations, including requirements, goals, and constraints, given to a technical system." This diversity underscores the importance of clear intent translation.
Although IBN offers significant potential, a major challenge lies in accurately and dynamically translating high-level user intents into precise RAN instructions. For IBN to work well and for networks to become fully autonomous, how accurately and reliably intent is translated is crucial. Understanding intents, which use natural language, requires intelligent translation. Incorrect translations can result in misconfigurations and severe network failures [6].
The advent of LLMs presents a powerful opportunity to overcome this critical translation gap. Some popular LLMs demonstrate excellent performance in processing text input, excelling in their ability to understand and generate natural language text [7, 8]. This makes them ideal as intelligent intermediaries capable of interpreting formalized intents and generating highly granular RAN configurations. Their advanced semantic understanding allows them to capture the subtle nuances and context inherent in complex service requirements, a significant leap beyond traditional rule-based or statistical methods.
The close loop mechanism of an IBN system, as defined in [2], comprises five components: Intent Profiling, Intent Translation, Intent Resolution, Intent Activation, and Intent Assurance. The interaction of these five components enables the IBN system to provide autonomous closed-loop operations. Intent Translation becomes the most critical and complex component. The translation process requires not only converting syntax but also understanding the meaning to accurately capture a user’s intent. Misinterpretations at this stage can lead to inefficient resource utilization, poor Quality of Service (QoS), or even network instability. The effectiveness of an intent-based RAN system is therefore fundamentally constrained by the accuracy and adaptability of its intent translation capability.
Existing studies on closed-loop network management, such as [9], translate intent using catalog databases. These approaches lack the advanced semantic understanding offered by LLMs. Separately, work in [10] uses Generative AI for intent management. However, it misses a crucial closed-loop mechanism, preventing continuous self-adaptation. The survey by [11] highlights key challenges in IBN, specifically concerning its goal to enhance automation and simplify network configurations. A major hurdle is the absence of a formal definition for "intent," primarily because various stakeholders hold differing views on its significance. Crucially, this underscores the importance of intent translation. It is the essential step for an IBN system to accurately interpret users’ high-level goals and requirements into actionable network policies.
This paper introduces an Intent-Based Wireless Network for RAN configuration that integrates LLMs as a core component to enhance intent translation. We elaborate on a structured prompt technique that leverages the cutting-edge semantic capabilities of LLMs. We demonstrate the system’s capability to dynamically optimize RAN parameters in an energy efficiency use case through a closed-loop mechanism. The manuscript is organized as follows. Section II discusses the proposed architecture. Section III discusses the results of experiment. Finally, Section IV addresses the lesson learned from our experiment and potential future works.
II Proposed Architecture
This paper presents an intent-based wireless network management system, as shown in Fig. 1. This system leverages cell configurations and cell measurement functions provided by the O-RAN experimental platform[12]. Crucially, the interface establishing the connection between our system and the RAN adheres to O-RAN specifications.
Our IBN system uses an agentic approach. A Strategist Agent, responsible for intent translation, receives the formalized intent declared by the consumer. This agent leverages a LLM to perform the translation, using the user prompt structure shown in Fig. 2.
Agentic LLMs serve as the intelligent engine that translate the structured intents and generate configuration strategy. This involves a multi-stage process. First, the agent analyzes the intent, understanding the target objects, the desired metric, the target condition, and the target value. Then agent then decomposes this objective into a series of actionable strategies. For planning and orchestration, the agentic LLM, orchestrated through a LangGraph flow[13], determines the necessary steps to achieve the intent. This may involve interacting with the RAN simulator via O-RAN O1 interface to retrieve relevant cell details and performance metrics. The "Strategist Agent," as shown in Fig. 1, plays a key role in formulating a configuration strategy based on the interpreted intent and potentially leveraging past successful attempts stored by the "History Analyzer Agent." The "Orchestrator Agent" then takes charge of executing these strategies by calling the appropriate configuration tools via the O-RAN O1 interface to configure the RAN simulator. While the focus of closed-loop monitoring might lean towards deterministic approaches, the initial configuration and optimization driven by the agentic LLM set the stage for efficient network operation.
II-A Formalized Intent Structure
Declaration of intent by owner is formalized as a JSON object, referencing 3GPP TS 28.312[14]. This formalized intent is then parsed to populate the Intent context, which serves as a structured input for the system. As depicted in Fig. 2, the Intent context primarily comprises two distinct sections: Target KPIs and Object Target. The Target KPIs section specifies the performance metrics to be monitored and optimized (e.g., energy efficiency), along with their desired conditions and target values. Conversely, the Object Target section identifies the specific network elements or scopes to which the intent applies (e.g., a particular cell or sub-network). For instance, if a user wants to ensure that the RAN energy efficiency for Cell 7 within SubNetwork 1 is optimized to be as high as possible, specifically targeting a value greater than 800,000 bits per Joule, an example intent might specify the Object Target as follows:
"objectInstance": "SubNetwork_1",
"ObjectTarget": ["Cell_7"]
And it declared alongside Target KPIs:
"targetName": "RANEnergyEfficiency",
"targetCondition": "IS_GREATER_THAN",
"targetValue": "800000",
"targetUnit": "bit/joule"
This comprehensive and structured representation of the intent serves as the initial input for our agentic LLM.
Figure 1: System architecture of an intent-based wireless network management system
II-B Prompt Techniques
We use one-shot prompting to effectively guide the LLM towards desired outputs. This technique provides a clear example of the expected output to ensure a consistent outcome. Specifically, we implement a five-section prompt structure to standardize input for LLM-driven intent translation. This structured approach is crucial for ensuring consistent and predictable LLM behavior, which is essential for the reliable and automated operation of an Intent-Based Network for RAN Management. Each section of the prompt is explicitly designed to provide comprehensive context and clear directives, allowing for precise control over the LLM’s decision-making process. The distinct purposes of these sections are detailed below:
  • Instruction: Each prompt begins with an Instruction, explicitly stating the task the LLM is expected to perform.
  • Intent: Contain the high-level objective or desired outcome from the system’s perspective.
  • Current Observation: Provides real-time or relevant operational data to give context about the current state of the network.
  • Configuration Constraints: This section is included to define any limitations for each RAN gNB, informing the LLM know what parameter and strategy are available to adjust.
  • Output Format: Specifies the exact structure in which the LLM’s response should be delivered. Defining structure output enables direct parsing for further computation processes.
Figure 2: User prompt structure for LLM agent
II-C Agentic Framework
Our system is built upon two distinct agents: the History Analyzer Agent and the Strategist Agent. The History Analyzer Agent is responsible for analyzing past strategies applied in the network and obtains real-time measurements directly from the RAN as input. Data about past strategy attempts are obtained from the Historical Strategy Attempts database. The History Analyzer Agent recommends the best strategy outcomes to the Strategist Agent. This processed data, along with a number of the most relevant previous strategy attempts, is then passed to the Strategist Agent for decision-making. Predefined strategy that contains data about target cell capabilities stored in the Strategy Templates database. Both agents perform their respective inference tasks calling the NVIDIA NIM inference API with detail inference as summarized in Table I. The underlying intelligence for these agents is configured within the Prompt Processor component, which utilizes the LangGraph library  [^1] https://github.com/langchain-ai/langgraph.
TABLE I: LLM Inference Parameters
Parameter Value
Model llama-3.1-70b-instruct
Temperature 0.2
Top P 0.7
Max Tokens 1024
Stream True
II-D O1 Configuration & Performance Management
To support the agentic architecture described in our Intent-Based Network for RAN Management, we leverage both Configuration Management (CM) and Performance Management (PM) over the O-RAN O1 interface to enable closed-loop network control. The agents continuously retrieve real-time network PM data which serve as critical input to assess the current operational state and evaluate the impact of previous configurations. For configuration, CM operations are performed using the NETCONF protocol. The agents first perform <get-config> operations to retrieve current RAN configuration parameters (e.g., transmit power levels). Using interpreted user intent and historical performance context, the Strategist Agent subsequently issues <edit-config> operations to apply optimized configurations to the RAN environment. This integration of CM and PM forms the backbone of the intent-based closed-loop control system.
This paper presents an intent-based wireless network management system, as shown in Fig. 1.
III Experimental Result
This section presents the results of our experimental analysis to validate the effectiveness of the proposed LLM-based RAN management approach. The experimental environment was established using the VIAVI Solutions AI RSG RAN Simulator, which was configured with the parameters listed in Table II.
TABLE II: Scenario Simulation
Parameter Value
Simulator VIAVI AI RSG
Version 2.4
Number of gNB 2
Number of UE(s) at gNB1 10
Number of UE(s) at gNB2 8
Mobility Model short range trajectory
Movement speed 0
III-A Single-Run Evaluation
[…]
Figure 3: Correlation between configured TxPower and PEE.EnergyEfficiency over Iteration(s) in a single-run test.
III-B Closed-Loop Precision Evaluation
Figure 4 illustrates the experimental results of a closed-loop precision evaluation for an intent-based system, focusing on the system’s performance across seven attempts with varying target values. The x-axis represents the number of attempts, ranging from 1st to 7th, while the y-axis indicates the number of iterations taken to achieve the target. An iteration indicates a cycle where the system performs a configuration, waits for the next PM sent by the target gNB, and completes the cycle upon receiving the PM. The first five attempts correspond to target value of 800. Then at 6th attempt we increase to 803 and finally is 805. The figures shows successful convergence with iteration counts of 3, 3, 2, 2, 2, and 4, respectively, reflecting the system’s ability to adapt to different precision levels.
The variation in iteration counts highlights the influence of the LLM responses on the system’s efficiency. Lower iteration numbers, such as 2, are observed when the LLM suggests more aggressive adjustments toward the target. In contrast, a higher iteration count of 4 at the sixth attempt suggests a more conservative adjustment process. The seventh attempt, marked by a "Fail" result with 7 iterations, indicates that the maximum allowable value of txPower to be adjusted was reached, leading the LLM to conclude a plateau. As a result, the system abandoned further optimization, as it was unable to fulfill the intent.
1st2nd3rd4th5th6th7th222333444777# of Attempt# of IterationSuccessFail
Figure 4: Closed-loop precision evaluation for an IBN System
III-C Intent Boundary Testing
Table III presents the results of the Intent Boundary Testing, which evaluates the system’s performance relative to the target PEE.EnergyEfficiency. The baseline value was established by recording initial simulation data using the VIAVI AI RSG prior to implementing any configurations. Subsequently, values both above and below this baseline were tested to assess system behavior. The results indicate that the system can process targets below the baseline, such as 750 kWh, by performing only monitoring and verification tasks, confirming that the system is already in a preferred state as defined by the intent. In contrast, when target PEE.EnergyEfficiency set above baseline threshold starting at 803 kWh, the system attempts to fulfill the target until it reaches a plateau, beyond which no further optimization is possible, leading to abandonment of the process. This behavior is also corroborated by the findings in Fig. 4.
TABLE III: Intent Boundary Testing for PEE.EnergyEfficiency
PEE.EnergyEfficiency (KWh) Description Result
765.4 Baseline -
>803 Above Baseline Threshold Failed
750 Below Baseline Threshold Passed
IV Lesson Learned and Future Works
This study presents a novel Intent-Based Network for RAN Management that harnesses LLMs to optimize the management of RAN configurations. Our primary contribution lies in developing a formalized prompt engineering technique that translates high-level user intents into a structured JSON format, encapsulating intent, network topology, KPIs, and configuration parameters, thus providing a standardized interface for LLMs to interpret and achieve the desired network objectives. We further introduce an agentic structure where two LLM agents collaboratively refine network configurations through iterative optimization in a close-loop manner. This iterative process ensures the intent is achieved through gradual parameter adjustments based on real-time feedback thereby preventing network disruptions caused by abrupt configuration changes.
However, during our preliminary evaluations, we identified two key limitations that critically impact the reliability and efficacy of LLM-driven network management: challenges related to contextual understanding and data accuracy.
IV-A Contextual Understanding and Domain Specificity
During the preliminary evaluation of our system, we observed that feeding comprehensive PM data directly to the LLM sometimes led to hallucinations or inconsistent interpretations in the generated configuration strategy. This highlights a critical lesson: while LLMs possess remarkable capabilities in natural language understanding, their ability to generate accurate and reliable responses within highly specialized domains, such as RAN configuration, is significantly enhanced by providing them with domain-specific ground truth.
This challenge highlights the need for future research to use a structured approach to serve data, explicitly detailing the relationships between entities. Such ground truth, ideally represented through formalisms like RDF (Resource Description Framework) or ontology, contains the precise nature of network entities and the fundamental properties of parameters (e.g., the exact relationship between a configuration parameter and a measurement metric). Representing and leveraging this complex domain knowledge before it is fed into the LLM is crucial for generating accurate, concise, and scientifically sound ’justification’ texts that the LLM can translate effectively.
IV-B The Impact of PM Data Inaccuracy
The operational efficacy of LLMs within a network management framework is predicated on its direct interaction with, and inherent assumption that, PM data represents the ’ground truth.’ A significant vulnerability arises from this unquestioning reliance, as our experiments revealed a critical aggregation issue. Although the RAN simulator transmits PM metrics at fixed intervals, the aggregated values often do not accurately reflect the network’s true state due to the various aggregation algorithms and configurable time windows.
When PM data is inaccurate, it sets off a series of consequences. A flawed initial understanding leads the LLM to make suboptimal decisions, which in turn generate new, equally unreliable PM data. This feedback loop of misinformation can rapidly destabilize the network, ultimately undermining the purpose of the automated management system.
Therefore, future research must address this challenge by exploring alternative data collection methodologies. One promising solution is to use PM metrics as a streaming data source without aggregation at the gNB to ensure all data is treated as a real-time reflection of the network state. Furthermore, integrating reinforcement learning (RL) algorithms can enhance system resilience by allowing the system to adapt to noisy or outlier data via iterative feedback and reward-based optimization. This will advance the development of robust and intelligent RAN management.
Acknowledgment
This work was supported by the National Science and Technology Council, Taiwan, under Contract numbers 111-2221-E-011-069-MY3, 112-2218-E-011-004, and 112-2218-E-011-006.